A boat leaves a dock and travels at the rate of 36 kilometers per hour. Twenty
minutes later a second boat leaves the same dock and travels the same route at 50 kilometers per
hour. How long will it take the second boat to be 5 kilometers behind the first boat?

Answer by greenestamps(13200)



The first boat travels at 36km/h for 20 minutes (1/3 hour) before the second boat starts, traveling (1/3)*36 = 12km.

In order for the second boat to become 5km behind the first, the second boat has to travel 12-5 = 7km farther than the first.

Related to the familiar formula

(distance) = (rate) times (time)

we have the corresponding formula

(difference in distances) = (difference in rates) times (time)

The difference in distances is 7km, and the difference in speeds is 50-36 = 14km/h, so

7 = (14)*t
t = 7/14 = 1/2

ANSWER: It will take the second boat 1/2 hour, or 30 minutes, to get to where it is 5km behind the first boat.